# Enhancing Multimodal Sentiment Analysis for Missing Modality through Self-Distillation and Unified Modality Cross-Attention

### v1

@@ I Introduction
Theoretically, textual information can be highly abstracted from audio and visual modalities. With advancements in Large Language Models (LLM), some studies [22, 23] have employed LLMs as decoders to address ASR tasks. For instance, the SLAM-ASR [22] projects WavLM’s [24] audio representations into the text space, and inputs them with prompt into an LLM to generate ASR results directly. Consequently, it is natural to infer that simulated text representations can be derived from the projected audio representations via LLMs, facilitating unified multimodal sentiment analysis.
@@
Our research proposes a Double-Flow Self-Distillation Framework, which includes the Modality Imagination Autoencoder (MIA) and Unified Modality Cross-Attention (UMCA). This unified model can address both complete and text-missing multimodal sentiment analysis scenarios by simply activating or deactivating the MIA during inference. The main contributions of our work are: (1) We design UMCA for comprehensive multimodal fusion to extract sentiment-related information. (2) Our method uses the LLM-based model Vicuna [25] to generate simulated text representations from WavLM’s hidden state projection when text modality is missing, with MIA complementing the missing information. (3) We train a unified multimodal sentiment analysis network using self-distillation combined with MIA activation and deactivation, while achieving distance-space alignment of complete and missing modality representations through MKD Loss and RNC Loss[26]. (4) Our method achieves optimal performance on the CMU-MOSEI [27] across multiple metrics for complete modality inference, and significantly improving performance in text-missing scenarios compared to other methods.
@@ II Method
In this section, we discuss our proposed Double-Flow Self-Distillation Framework for the multimodal sentiment analysis task. The overall architecture of our system is depicted in Fig. 1 and will be detailed in the subsections.
@@ III-D Overall Comparison
We compare our method with other high-performing approaches that have achieved excellent performance or studied modality missing. In detail, the models listed in Table II explore robustness against missing modalities: GMC uses contrastive learning to align representations, GCN employs graph neural networks, GMD reduces modality dependence via Gradient-guided Decoupling, MMIN reconstructs missing modalities, and MInD separates modality-invariant information and modality-specific information. We referenced the experimental results of these models obtained from articles on the GMD method. The results are shown in Table II.
@@ IV Conclusions
Our work addresses the challenge of the high cost of acquiring text modality and the strong reliance of multimodal sentiment analysis on it. Specifically, we designed the Double-Flow Self-Distillation Framework with Unified Modality Cross-Attention as the main network structure, combined with Modality Imagination Autoencoder to simulate missing text modality. Additionally, we developed a series of loss functions to enhance the model’s performance in multimodal sentiment analysis, ensuring robustness both when all modalities are present and when the text modality is missing.

### v2

@@ I Introduction
Theoretically, textual information can be highly abstracted from audio and visual modalities. With advancements in Large Language Models (LLM), some studies [22, 23] have employed LLMs as decoders to address ASR tasks. For instance, SLAM-ASR [22] projects audio representations and inputs them with prompts into an LLM to generate ASR results. Consequently, it is natural to infer that simulated text representations can be derived from the projected audio representations via LLMs, facilitating unified multimodal sentiment analysis.
@@
Our research proposes a Double-Flow Self-Distillation Framework, which includes the Modality Imagination Autoencoder (MIA) and Unified Modality Cross-Attention (UMCA). This unified model can address both complete and text-missing multimodal sentiment analysis scenarios by simply activating or deactivating the MIA during inference. The main contributions of our work are: (1) We design UMCA for comprehensive multimodal fusion to extract sentiment-related information. (2) Our method uses the LLM-based model Vicuna [24] to generate simulated text representations from audio representations’ projection when text modality is missing, with MIA complementing the missing information. (3) We train a unified multimodal sentiment analysis network using self-distillation combined with MIA activation and deactivation while achieving distance-space alignment of complete and missing modality representations through MKD Loss and RNC Loss[25]. (4) Our method achieves optimal performance on the CMU-MOSEI [26] across multiple metrics for complete modality inference, and significantly improving performance in text-missing scenarios compared to other methods.
@@ II Method
In this section, we discuss our proposed Double-Flow Self-Distillation Framework for the multimodal sentiment analysis. The overall architecture of our system is depicted in Fig. 1.
@@ III-D Overall Comparison
We compare our method with other high-performing approaches that have achieved excellent performance or studied modality missing. The approaches listed in Table II explore robustness against missing modalities using various methods: contrastive learning (GMC), graph neural networks (GCN), modality dependence reduction via gradient-guided decoupling (GMD), missing modality reconstruction (MMIN), and modality information separation (MInD), et al. We referenced the experimental results of these models obtained from articles on the GMD method. The results are shown in Table II.
@@ IV Conclusions
This work addresses the challenge of high cost and heavy reliance on the text modality. Specifically, we designed the Double-Flow Self-Distillation Framework with Unified Modality Cross-Attention as the main network structure, combined with Modality Imagination Autoencoder to simulate missing text modality. Additionally, we developed a series of loss functions to enhance the model’s performance in multimodal sentiment analysis, ensuring robustness both when all modalities are present and when the text modality is missing.
V Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62171427.